OpenStudy (anonymous):

3/4 = 15/? I got 20, am I right ?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Yay aha thank you (:

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

Hey, I have a few more I'm not too sure on, would you please help me ?

OpenStudy (anonymous):

Okay sure!

OpenStudy (anonymous):

It covers a bit of everything, its not just fractions.

OpenStudy (anonymous):

round 495 to the nearest 10. 500 ?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

15 % of $321.00 $48.60 ?

OpenStudy (anonymous):

Wouldn't it be $48.15?

OpenStudy (nubeer):

yeah it would be 48.15

OpenStudy (anonymous):

So how did you get that ?

OpenStudy (nubeer):

15% means 15/100 \[\frac{ 15 }{ 100 } \times 321\]

OpenStudy (anonymous):

awk, ive pressed the wrong button somewhere aha, thank you (:

OpenStudy (nubeer):

lol no problem

OpenStudy (anonymous):

-5 x -3 = ? I got 15

OpenStudy (usukidoll):

ya

OpenStudy (nubeer):

right..

OpenStudy (nubeer):

if that x is multiply

OpenStudy (anonymous):

yeah it is aha

OpenStudy (anonymous):

I think thats all for this section aha. Thank you both (:

OpenStudy (nubeer):

you're welcome. :)